OpenStudy (anonymous):

find an equation of the line having the given slope and containing the given point m=-2, (1,8)

OpenStudy (anonymous):

y=mx+c y2-8=-2*(x2-1) y2=-2x2+10

OpenStudy (anonymous):

the equation of a line is y=mx+c. since m=-2,y=-2x+c then substitute x=1,y=8(since the point (x,y)=(1,8)in y=-2x+c 8=-2(1)+c 8=-2+c 10=c therefore the equation is y=-2x+10